calculate the unit step response of a system with the difference equation

y[n]+1.5y[n-2]=x[n]-x[n-1], y[-1]=2 ,y[-2]=1
Answer #1

E The difference equation given is as follows :- y [n] + 1-5 y[n-2] = x[n]-x [n-1] => z fy [n]+ 1-5 y[n-2)} = 24 x [n] – x[n->> Y(3) +1.54 z? Y(z)+ z yE] + y(-3]} = X(z) – +2X(z) => (1+1.52-2) Y(z) +(1.5z+1)(2) +(1:5)(1) = X(3) (1-21) Since, we arenEO » X(z) = Eur] zen >> X(z) = 1+ 2++32+..... » X(z) - t (Applying geometric (1-z-1) progression to infinite We took commonputting the value and rearranging eq@, 2247 Ye3 = cos(s) or ( = Y(z) = -0.51 t 1- 171522) 449 1 1.2247 1+1.522 taking inverse